a piggy bank contains $5.85 in nickels and quaters. If there are 65 coins in all, how many of each type coin is there?
Answer by checkley75(3666)

N+Q=65 OR N=65-Q
.05N+.25Q=5.85
.05(65-Q)+.25Q=5.85
3.25-.05Q+.25Q=5.85
.20Q=5.85-3.25
.20Q=2.60
Q=2.60/.20
Q=13 QUARTERS.
65-13=52 NICKELS.
pROOF:
.05*52+.25*13=5.85
2.60+3.25=5.85
5.85=5.85